If you had a liquid culture of Staphylococcus aureus and transferred 3 ml to test tube “R”, which contains 6 ml of nutrient broth, what would be the final dilution (in simplest terms) of the culture in test tube “R"? From the culture in “R” make a 1:6 dilution in tube “S” (this is not the final or overall dilution). Now, determine what is the final or overall dilution of the culture in tube “S”? SHOW ALL YOUR WORK
Answer #1

Answer:

Given that 3 ml of Staphylococcus aureus culture is added to 6 ml nutrient broth.

Dilution

= sample volume/total volume

=3/(3+6)

=3/9

=1/3

Thus, final dilution of culture in test tube "R" is 1/3.

Again, from tube "R" a 1:6 dilution is made to form"S".

Dilution in tube "S"

=1/3 x 1/6

=1/18

Hence, final dilution in tube "S" is 1/18.